40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 17 16:44:04 2015, Koji, Update, LSC, Delay line installed again 
    Reply  Tue Feb 17 19:49:51 2015, Koji, Update, LSC, Delay line un-installed again 
       Reply  Mon Feb 23 21:57:13 2015, Koji, Update, LSC, Delay line installed again (experiment, round 1) 
          Reply  Tue Feb 24 01:57:01 2015, Jenne, Update, LSC, Modelled effect of relative modulation phase ChangeInAngle_PRCLvsMICH.pngREFL33_MICH_SensMat_vsModulationOrder_vsModulationPhase.pngREFL33_SensMat_vsModulationPhase.pngREFL33_PRCL_nominalVsLoweredGamma.pngREFL33_MICH_nominalVsLoweredGamma.png
             Reply  Thu Feb 26 00:20:54 2015, ericq, Update, LSC, Modelled effect of relative modulation phase 33budget_canceled.png33budget_nominal.png2015-02-ModPhase.zip
                Reply  Thu Feb 26 01:51:39 2015, ericq, Update, LSC, Sideband HOMs C1_HOMcurves_PR.pngC1_HOMcurves_Y.pngC1_HOMcurves_X.pngC1_HOMlist.zip
                Reply  Thu Feb 26 01:53:35 2015, Koji, Update, LSC, Modelled effect of relative modulation phase 
       Reply  Tue Feb 24 23:39:16 2015, Jenne, Update, LSC, Delay line re-installed again 
          Reply  Wed Feb 25 04:59:47 2015, Jenne, Update, LSC, Delay line re-installed, measurements round 2 SensMat_PRMI_24Feb2015.pdf
             Reply  Wed Feb 25 12:16:27 2015, Koji, Update, LSC, Delay line re-installed, measurements round 2 
                Reply  Wed Feb 25 14:18:28 2015, rana, Update, LSC, Delay line re-installed, measurements round 2 
          Reply  Wed Feb 25 14:51:13 2015, Jenne, Update, LSC, Delay line un-installed again 
             Reply  Wed Feb 25 20:00:39 2015, Jenne, Update, LSC, Delay line un-installed again - sensing matrix comparisons 
                Reply  Wed Feb 25 23:48:57 2015, rana, Update, LSC, some thoughts 
                   Reply  Thu Feb 26 20:02:46 2015, Jenne, Update, LSC, re: some thoughts 
Message ID: 11064     Entry time: Wed Feb 25 04:59:47 2015     In reply to: 11062     Reply to this: 11066
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Delay line re-installed, measurements round 2 

[Jenne, EricQ, Rana]

We spent this evening measuring and thinking about our 3f signals, and the effect of the modulation cancellation. 

I reinstalled the delay line box, and reduced the modulation depth of the 55MHz signal, so that we are in the state of modulation cancellation - there should be almost no power at 33MHz injected into the vacuum.  I carefully tuned the demod phase of the REFL 11, 33 and 55 MHz PDs, and locked the PRMI on REFL55 I&Q.  The signal in REFL 165 was very tiny, so as best as I could tell, the demod phase that Koji found last week was correct. 

Here is a little record of what the demodulation phases should be, for the "nominal" and "cancellation" configurations, so that we don't have to continually use the time machine.

  "Nominal" configuration 3f modulation cancellation
REFL 11 20 deg 76.0 deg (tuned to nearest 0.5 deg)
REFL 33 142.2 deg 120.3 deg (tuned to nearest 0.1 deg)
REFL 55 19 deg 173.0 deg (tuned to nearest 0.5 deg)
REFL 165 -172 deg 18 deg (same number as last week)
AS 55 -166.1 deg -111.1 deg (same number as last week)

Also, here is the locking recipe for REFL55 I&Q in the cancellation configuration:

PRMI, 3f cancellation, REFL55 I&Q MICH PRCL
Input Matrix acquire with -2*REFL55Q, then put matrix to -15*REFL55Q -12*REFL55I
Gain 3.0 -0.1
DoF trigger POP22I; 50 up, 0.5 down POP22I; 50 up, 0.5 down
FMs FM 4, 5 on FM 4, 5 on
FM trigger FM 1, 2, 3, 6, 9; 50 up, 0.5 down, 5 second wait FM 1, 2, 6, 9; 50 up, 0.5 down, 1 second wait
Normalization none none
Output matrix -1*ITMX, +1*ITMY 1*PRM

With this setup, we measured the sensing matrix.  MICH had an excitation at 370.123 Hz with 8,000 counts to -ITMX+ITMY (this is about 0.3nm for each ITM), and PRCL had an excitation at 404.123 Hz with 50 counts to PRM.  For tonight, here is a PDF of the peaks in DTT.  The data is saved in /users/Templates/LSC_error_spectra/SensMat_PRMI_24Feb2015.xml. 

SensMat_PRMI_24Feb2015.pdf


Rana has proposed to us an idea for why the REFL 33 signal should be dominated by the 55*22 contribution, rather than -11*22.  Eric is in the process of checking this out with a Mist model to see if it makes sense. Here's the gist:

Our Schnupp asymmetry is small (3.9cm, IIRC), so the transmission of the 11MHz signal out the dark port is small.  This means that the finesse of the PRC for 11MHz isn't so huge.  On the other hand, since 55MHz is a higher frequency, the transmission out the dark port is larger and is a closer match to the PRM transmission, so the finesse of the PRC for 55MHz is higher. 

Since the magnitudes of the fields at the reflection port are not changing significantly, our PDH signals are being created by the relative phase of something which is anti-resonant (ex. carrier or 22MHz for sideband lock) vs. something which is resonant (ex. 11MHz or 55MHz).  Since the finesse of the 55MHz signal is larger, the accumulated phase change is greater, so we expect a larger slope to our PDH signals that involve 55MHz as compared to those that use 11MHz.

If we are comparing the contributions between -11*22 and 55*22, they both include the anti-resonant 22MHz. So, the difference in the signal strengths comes directly from the difference in phase accumulation due to the variation in the dark port transmission.

EricQ had a thought, and while I have enough awake brain cells to report the thought, we're going to have to revisit it when more of our brains are awake.  In either case, the transmission through the dark port is small compared to the transmission of the ITMs, so why don't the ITMs dominate the finesse calculation, and thus are the 11MHz and 55MHz really getting that much of a difference in finesse?  To be checked out.

Attachment 1: SensMat_PRMI_24Feb2015.pdf  15 kB  | Hide | Hide all
SensMat_PRMI_24Feb2015.pdf
ELOG V3.1.3-